OpenStudy (anonymous):

Find a quadratic function associated with the vertex (h,k)=(2,1) and point (x,y)=(6,7) that the parabola passes through.

OpenStudy (jdoe0001):

you've covered parabolas I assume?

OpenStudy (anonymous):

Yes.

OpenStudy (campbell_st):

well I use the standard form (x - h)^2 = 4a(y - k) (h, k) is the vertex so substitute those values... then substitute the point to find the value of a. then you have the equation... to get it to vertex form is simply a case of making y the subject

OpenStudy (anonymous):

Okay, thanks!

OpenStudy (jdoe0001):

ok so we know the vertex is at (2,1) \(\bf y=a(x-{\color{brown}{ h}})+{\color{blue}{ k}}\qquad vertex\ ({\color{brown}{ 2}},{\color{blue}{ 1}})\implies y=a(x-{\color{brown}{ 2}})+{\color{blue}{ 1}}\)

OpenStudy (jdoe0001):

well.. using campbell_st's will work the same :)

OpenStudy (jdoe0001):

and use the vertex template, with the vertex coordinates and since the line passes through (6,7) meaning x =6, y =7 then just plug that in the vertex form and solve for "a" to find that factor then plug it back in the equation